OpenStudy (anonymous):

The values of x satisfying the equation 8x^3/2n - 8x^-3/2n = 63 are?

OpenStudy (anonymous):

@Callisto @experimentX @badi @dpaInc @FoolForMath @JamesJ @jerwyn_gayo

OpenStudy (anonymous):

The answer should be 2^2n and 1/(2^2n)

OpenStudy (experimentx):

let, x^3 = u

OpenStudy (anonymous):

ok

OpenStudy (experimentx):

x^-3 = 1/u and solve quadratic equation

OpenStudy (anonymous):

It is not cube it is x^(3/2n)

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

oh .. i believe both 3 is 3n??

OpenStudy (pfenn1):

Is this your equation?\[8x^{\frac{3}{2n}} - 8x^{-\frac{3}{2n}} = 63 \]

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

So the 8 is not affected by the exponent, only the x?

OpenStudy (callisto):

\[8x^{\frac{3}{2n}} - 8x^{-\frac{3}{2n}} = 63\]\[x^{\frac{3}{2n}}(8x^{\frac{3}{2n}} - 8x^{-\frac{3}{2n}}) = 63(x^{\frac{3}{2n}})\]\[8x^{2(\frac{3}{2n})} - 8 = 63x^{\frac{3}{2n}}\]\[8x^{2(\frac{3}{2n})} -63x^{\frac{3}{2n}} - 8 = 0\]\[(8x^{\frac{3}{2n}}+1)(x^{\frac{3}{2n}}-8)=0\]\[(8x^{\frac{3}{2n}}+1)=0 \ or\ (x^{\frac{3}{2n}}-8)=0\]\[x^{\frac{3}{2n}} = -0.125 \ or \ x^{\frac{3}{2n}} =8\]

OpenStudy (callisto):

That's weird :|

OpenStudy (anonymous):